To calibrate your calorimeter cup, you first put 47 mL of cold water in the cup, and measure its temperature to be 25.9 °C. YI don't know why the last two are wrong.
Answer #1

The heat of the calorimeter is calculated:

q cal = 2222 - 2604 = - 382 J

Ccal is calculated:

Ccal = 382 / (37.2 - 25.9) = 33.81 J / ° C

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.